# Convolution goes higher-order: a biologically inspired mechanism empowers image classification

### v1

@@ A.3 Architectures and Training Procedure: MNIST, FashionMNIST, CIFAR-10, CIFAR-100 & Imagenette
A.3.1 Architectures and Parameter Counts: Imagenette
@@
A.3.2 Training Setup: Imagenette
@@
A.4 Comparison with VOneNet
We compared the performance of our HoResNet-18 architecture with a ResNet-18 adapted VOneNet (Dapello et al., 2020), a biologically-inspired model designed to better capture the properties of the visual cortex. Our HoResNet-18 model achieved superior accuracy on the Imagenette dataset (89.30%) compared to the VOneNet model (88.02%), while utilizing fewer parameters (11,168,382 vs. 14,445,322).
@@ A.4 Comparison with VOneNet
A.5 Tied-Weights Issue
@@
A.6 Sensitivity to Image Statistics
More detailed results are presented in Table 3 and 4 (in bold largest decrease in performance across models).
@@ A.6 Sensitivity to Image Statistics
A.7 Neural Representations: Representational Similarity Analysis
To gain deeper insight into how our Higher-order Convolutional layer (HoConv) processes information differently from a standard Convolutional (Conv) layer, we employed Representational Similarity Analysis (RSA) (Kriegeskorte et al., 2008) on the CIFAR-10 dataset. We analyzed activations from 100 test images (10 per class) averaged across 50 model realizations to mitigate the effects of random initialization (Ding et al., 2021). Our findings reveal distinct representational geometries between the two models. In Figure 8 we consider an alternative distance (hellinger) between the two RDMs, based on the consideration of the fact that values are between 0 and 1.
@@ A.7 Neural Representations: Representational Similarity Analysis
A.7.1 Order-wise Representational Differences
@@
A.7.2 Layer-wise Representational Differences
@@
A.7.3 Correlation between RDMs across Layers

### v2

@@ A.9 Architectures and Training Procedure: MNIST, FashionMNIST, CIFAR-10, CIFAR-100 & Imagenette
A.9.1 Architectures and Parameter Counts: Imagenette
@@
A.9.2 Training Setup: Imagenette
@@
A.10 Comparison with VOneNet
We compared the performance of our HoResNet-18 architecture with a ResNet-18 adapted VOneNet [45], a biologically-inspired model designed to better capture the properties of the visual cortex. Our HoResNet-18 model achieved superior accuracy on the Imagenette dataset (89.30%) compared to the VOneNet model (88.02%), while utilizing fewer parameters (11,168,382 vs. 14,445,322).
@@ A.10 Comparison with VOneNet
A.11 Tied-Weights Issue
@@
A.12 Sensitivity to Image Statistics
More detailed results are presented in Table 9 and 10 (in bold largest decrease in performance across models).
@@ A.12 Sensitivity to Image Statistics
A.13 Neural Representations: Representational Similarity Analysis
To gain deeper insight into how our Higher-order Convolutional layer (HoConv) processes information differently from a standard Convolutional (Conv) layer, we employed Representational Similarity Analysis (RSA) [17] on the CIFAR-10 dataset. We analyzed activations from 100 test images (10 per class) averaged across 50 model realizations to mitigate the effects of random initialization [47]. Our findings reveal distinct representational geometries between the two models. In Figure 8 we consider an alternative distance (hellinger) between the two RDMs, based on the consideration of the fact that values are between 0 and 1.
@@ A.13 Neural Representations: Representational Similarity Analysis
A.13.1 Order-wise Representational Differences
@@
A.13.2 Layer-wise Representational Differences
@@
A.13.3 Correlation between RDMs across Layers